OpenStudy (anonymous):

Is 170 % greater than the square root of 2.89

OpenStudy (greencat):

170%=170/100

OpenStudy (anonymous):

So it is not

OpenStudy (greencat):

170/100=1.7

OpenStudy (greencat):

\[\sqrt{2.89}=1.7\]

OpenStudy (anonymous):

ohhh

OpenStudy (greencat):

It is not greater, it is equal. So, basically, not greater.

OpenStudy (anonymous):

I have 1 more

OpenStudy (greencat):

Also, I advise that you use a calculator; these are simple functions.

OpenStudy (greencat):

I'll help you.

OpenStudy (anonymous):

Is 5 1/6 greater than the square root of 27

OpenStudy (jdoe0001):

*cough* "Also, I advise that you use a calculator; these are simple functions."

OpenStudy (greencat):

http://prntscr.com/5oz3yu ^The number is larger than 5 1/6.

OpenStudy (greencat):

@jdoe0001 What? Did I say anything wrong or mean? I'm sorry if I did.

OpenStudy (jdoe0001):

hmmm nope.... I was just reposting since nyprincess2001 seems she isn't reading what you posted :)

OpenStudy (greencat):

Oh ok.